The following data describe the distribution of grades out of 20, obtained by 20 students during a quiz of Statistics: 6; 7; 5; 7; 7; 8; 7; 6; 9; 7; 4; 10; 6; 8; 8; 9; 5; 6; 4; 8.

Calculate the mean.


1
Expert's answer
2021-05-07T09:00:02-0400

Mean =summation of X/total number (N)

Summation of X = 6+7+5+7+7+8+7+6+9+7+4+10+6+8+8+9+5+6+4+8=137

Total number=20

137/20=6.85

The mean is 6.85